OpenStudy (anonymous):

A company sells cardboard that is eight and thee-fourths millimeters thick. Write an equation to solve for the number of sheets of cardboard, c, in a pile that is one hundred fifty-seven and one-half millimeters thick. What must be done to both sides to solve the equation?

OpenStudy (texaschic101):

8 3/4c = 157 1/2 -- turn to improper fractions 35/4c = 315/2 -- divide both sides by 35/4 c = (315/2) / (35/4) c = 315/2 * 4/35 c = 630/35 c = 18 There are 18 sheets.

OpenStudy (anonymous):

thank you :) word problems like this confuse me

OpenStudy (texaschic101):

you would have to divide both sides by 35/4 (which is the improper fraction form of 8 3/4)

OpenStudy (texaschic101):

glad to help....word problems can be confusing sometimes